A fair 6 sided die is rolled twice, determine the probability of obtaining the sum of the numbers less than 5. 


1
Expert's answer
2022-01-25T04:39:57-0500

The total number of outcomes after rolling a dice twice is "6^2=36".

Pairs of numbers that satisfies the condition are: (1,1) (2,2) (1,2) (1,3) (2,1) (3,1)

6 pairs in total

So, the required probability is "{\frac 6 {36}}={\frac 1 6}"